# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs

### v6

@@ Abstract
LLM-as-a-Judge presents a promising alternative to human evaluators across various tasks, but inherent biases, especially position bias — a tendency to favor solutions based on their position in the prompt — have compromised its effectiveness. Our study introduces a systematic framework to examine position bias in pairwise comparisons, focusing on repetition stability, position consistency, and preference fairness. This research significantly contributes to the field by introducing new concepts for understanding position bias and providing a multi-dimensional framework for evaluations. We conducted experiments with 12 LLM judges across MTBench and DevBench, covering 22 tasks and approximately 40 solution-generating models — candidates, resulting in over 100,000 evaluation instances. Our findings confirm that position bias in capable LLM judges is not due to random chances, along with notable variations observed across judges and tasks. Moreover, position bias is weakly influenced by the length of prompt components but significantly impacted by the quality gap between solutions. These insights can help optimize judge model selections, improve benchmark design, and inform future research on debiasing strategies, ultimately enhancing the reliability of LLM judges.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating large language models (LLMs) has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations (Zeng et al., 2023; Karpinska et al., 2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments(Zheng et al., 2024b; Li et al., 2024; Zhu et al., 2023b), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1. This bias is prevalent across all categories of LLM-as-a-Judge, including pointwise, listwise, and pairwise evaluations (Qin et al., 2024), for any score-based and relation-based settings (Li et al., 2023d). Consequently, the reliability of LLM judges becomes questionable when they exhibit position bias. However, addressing this issue is highly complicated due to the varying performance of LLM judges across different tasks (Khan et al., 2024; Chua et al., 2024). Moreover, the limited understanding of the bias itself has exacerbated the problem.
Prior studies on position bias implicitly assume that this bias is not due to random variations. In other words, these studies assume that LLM judges produce consistent results across repeated trials. However, without validating this assumption, evaluations may become less reliable. As shown in Fig. 1, the occurrence of position bias can be observed by chance even for the same set of tasks and candidate models. Thus, it is unclear whether the observed position bias actually arises from the position of the prompt components or is merely due to random chance. Therefore, when exploring the position bias of LLM judges, it is necessary for the judges to be stable across repetitions to yield nontrivial results.
Figure 1: Demonstration of position bias and repetition stability of LLM-as-a-Judge. When the positions of candidate model solutions are swapped in the prompt, position bias is exhibited if the LLM judge tends to favor the same position (e.g., always choosing {B}) rather than the same candidate model solution. We observe repetition instability if the same LLM judge gives different judgments across repetitions. Only when an LLM judge has high repetition stability can we validate that its position bias is not due to random variations.
Another key aspect of position bias, preference fairness, has also been under-explored by previous works. Most studies have primarily focused on position consistency to measure how frequently position bias occurs across a set of evaluation instances. However, the direction of biased preference (i.e., primacy or recency) is also insightful. Fig. 2 illustrates cases where a pair of judgments with swapped prompts become position consistent, inconsistent, primacy-preferred, or recency-preferred. We measure preference fairness by analyzing the distribution of these preference directions across a set of evaluations. Specifically, judgments are considered to have a fair preference if they are position consistent or if the preference directions are evenly distributed. In contrast, an LLM judge is considered to have an unfair preference if it consistently favors candidate model solution that appear either first or last in the prompt. Both fair and unfair judge models can provide valuable insights into model properties, applications, and strategies for improvement.
Figure 2: Demonstration of position consistent/inconsistent, primacy-preferred, and recency-preferred judgment pairs in pairwise comparative assessment, where the task is to choose the better candidate solution. A judgment is consistent if it selects the same winner across different prompt positions. Primacy/recency indicates a preference for the first/last candidate solution, respectively. Position inconsistent judgments suggest position bias and can be classified as primacy-preferred or recency-preferred.
To enhance the understanding of position bias and LLM-as-a-Judge, we propose a comprehensive framework designed to dissect and analyze position bias within the context of pairwise comparative assessment. We focus on this context because LLM judges demonstrate superior performance in making pairwise comparisons over other settings like pairwise scoring (Zheng et al., 2024b; Liusie et al., 2024), making it a practically preferable choice for investigations.
Our contribution can be summarized as following: (1) introducing a systematic framework for evaluating LLM-as-a-Judge’s position bias in terms of Repetition Stability, Position Consistency, and Preference Fairness; (2) identifying Judge-level, Candidate-level, and Task-level factors that affect position bias; and (3) providing insights into future benchmark design and paving the way for effective mitigation strategies for position bias.
To better understand position bias and the effectiveness of LLMs as evaluators, we have developed a detailed framework specifically for analyzing position bias in scenarios where two options are compared directly, known as pairwise comparative assessment. We choose this approach because LLMs tend to perform better in these head-to-head comparisons than in other formats like scoring each option separately (Zheng et al., 2024b; Liusie et al., 2024). This makes pairwise comparisons a practical focus for our research.
Our contributions are threefold: (1) We introduce a systematic framework to analyze how LLMs judge by focusing on Repetition Stability, Position Consistency, and Preference Fairness. (2) We explore factors at different levels—Judge, Candidate, and Task—that influence position bias. (3) We offer insights that will help in designing better benchmarks and developing strategies to reduce position bias in LLM evaluatio
2 Methods & Defintions
LLM-as-a-Judge can be classified into two categories: score-based and relation-based assessments. under OuFirst, the pairwise paradigm is a popular and classic scenario that has influenced a wide range of fields (Qin et al., 2024). Second, as noted by Liu et al. (2024), LLM and human evaluations are more aligned in the context of pairwise comparison than any of the score-based assessments. Additionally, numerous studies have shown that pairwise comparative assessment outperforms other judging settings in terms of position consistency(Zheng et al., 2024b; Liusie et al., 2024). Finally, pairwise comparisons can be extended to more complicated frameworks of relation-based assessment, such as listwise comparisons, through some advanced ranking algorithms (Qin et al., 2024; Liu et al., 2024). Therefore, analyzing position bias in this context can provide a deeper understanding of its fundamental nature.
In the context of pairwise comparative assessments, LLM judges are prompted to choose the better candidate solution to the task question. To accommodate the possibility of a tie, various option modes are introduced. Two-Option mode restricts judges to choosing the better solution between two options, labeled {A} for the first and {B} for the second as shown in Fig. 1. Three-Option mode adds an additional choice, {C}, allowing judges to indicate a tie if neither solution is preferable. Four-Option mode further expands the choices, with {C} representing a both good tie and {D} a both bad tie. In our experiments, we employed Three-Option mode for MTBench (Zheng et al., 2024b) and Two-Option mode for DevBench (Li et al., 2024), aligning with their respective original frameworks. These option modes are explicitly specified in the system prompts to clearly guide the decision-making process of the LLM judges.
@@ 2 Methods & Defintions
To set up our experiment, as shown in Fig. 1, the system prompt, option mode, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the winning solution. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question. Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair. As shown in Fig. 2, if the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent with a fair preference. Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy or recency.
To systematically investigate position bias in pairwise comparative judging scenarios, we evaluate LLM judges using three key metrics: Repetition Stability ($RS$), Position Consistency ($PC$), and Preference Fairness ($PF$). While $PC$ has been widely adopted in prior research, we introduce $RS$ to validate the reliability of LLM judges and $PF$ to provide a more comprehensive assessment. Further analysis examines the impact of judge-level, candidate-level, and task-level factors on position bias. Therefore, for each judge, candidate, and task, we calculate $RS$, $PC$, and $PF$ to thoroughly explore the variance in position bias and its impacting factors.
@@ 2.1 Evaluation Metrics
We validate the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows.
@@
Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times. It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or merely random variations. We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset. This metric is formalized as
$RC=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{\forall k\in S}\left\{|C_{k}^{j}|\right\},\quad S=\{A,B,C,...\}$ (1)
@@
where $S$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries. The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
@@
Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is swapped. It is calculated as the ratio of consistent evaluation pairs to the total number of valid evaluations, where a pair is deemed consistent if the judge model prefers the same winning solution. Formally, it is calculated as
$PC=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})\in V\}},V=\{(A,B),(B,A),(C,C),...\}$ (2)
@@
where $(C_{\texttt{original}}^{j},C_{\texttt{swapped}}^{j})$ denotes the judgment pair for the $j^{th}$ query, $n$ represents the number of prompt pairs, and $V$ is the set of choice pairs that correspond to position consistency. An example of such pairs of choices can be found in Fig. 2. This formula provides a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
Preference Fairness ($PF$) is another crucial characteristic of position bias measuring the extent to which judge models favor certain solution positions. In the pairwise comparative setup, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second position” used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 2.
@@
Position consistent judgments are fair in preference. The measurements become more complex for position inconsistent cases. Previous studies proposed two common ways to measure the preference fairness. A straightforward way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023b). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
@@
Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach is problematic because it overlooks the fact that “position consistent judgments are also position fair,” which can lead to overly penalizing highly consistent LLM-judges. For instance, if only one out of a hundred pairs is inconsistent, this method would classify it as 100% primacy-preferred or recency-preferred. However, this outcome should be distinguished from cases where all one hundred pairs are inconsistent and biased toward one position, which would also result in 100% preference unfairness.
@@
$PF=\frac{PF_{raw}-S^{-}_{min}}{S^{+}_{max}-S^{-}_{min}}\times 2-1,PF_{raw}=(rcn\times irr)-(pcn\times ipr)$ (3)
where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively. This min-max scaling technique ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero. Compared to prior methods, the use of $rcn$, $pcn$ weighted by $irr$ and $ipr$, and scaling across judges and tasks implicitly accounts for the position consistent judgments. Another advantage of this approach is that it quantifies preference fairness with a single score, rather than the two separate scores used in previous studies. The $PF$ score is interpreted as follows:
@@
To summarize, our proposed preference fairness score is superior to prior measurements due to its ability to provide a single, comprehensive score that accounts for all evaluation instances, while maintaining its sensitivity across datasets, even when the number of questions and instances varies for each task.

### v7

# Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge

@@ Abstract
LLM-as-a-Judge has emerged as a promising alternative to human evaluators across various tasks, yet inherent biases—particularly position bias, the tendency to favor solutions based on their position within the prompt—compromise its reliability. This study investigates position bias in LLM judges across pairwise and list-wise comparison settings, introducing three metrics: repetition stability, position consistency, and preference fairness. Our experiments, involving 12 LLM judges across MTBench and DevBench with 22 tasks and approximately 40 solution-generating models, result in over 100,000 evaluation instances. We identify Judge-Level, Candidate-Level, and Task-Level factors contributing to bias. The findings confirm that position bias is not due to random chance and varies significantly across judges and tasks. While position bias is weakly influenced by the length of prompt components, it is strongly affected by the quality gap between solutions. Our agreement and disagreement analysis among judges further provides insights into the distribution of judging difficulty across the dataset, and highlights the potential for dataset modifications.
Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge
Lin Shi, Chiyu Ma, Wenhua Liang, Weicheng Ma, Soroush Vosoughi Dartmouth College {lin.shi.26, chiyu.ma.gr, wendy.liang.25, weicheng.ma.gr, soroush.vosoughi}@dartmouth.edu
@@ 1 Introduction
Figure 1: Overview of our experiment settings: (a) Position bias is observed when LLM judges consistently favor a specific position rather than evaluating the content, with repeated trials ensuring the deviations are not due to random variations. (b) Preference fairness is defined and measured through the distribution of choice pairs to assess the fairness of judgments. (c) The settings are extended from pairwise comparisons to list-wise comparisons, involving evaluations of more than two candidate models.
@@
In recent years, Large Language Models (LLMs) have emerged as evolutionary technologies, gathering global interest and stimulating substantial research into their applications. Evaluating LLMs has received increasing attention due to their advancing capabilities across diverse fields. While human assessment is considered the gold standard for aligning with human preferences, it lacks scalability and reproducibility in extensive evaluations Zeng et al. (2023); Karpinska et al. (2021). To automate evaluations and reduce reliance on costly human evaluators, the LLM-as-a-Judge methodology emerged as a promising alternative across various tasks. Despite a high level of agreement with human judgments Zheng et al. (2024b); Li et al. (2024a); Zhu et al. (2023), inherent biases, especially position bias, have undermined the accuracy, fairness, and reliability of these LLM evaluators.
Position bias refers to the tendency of LLM judges to favor certain positions within prompt components rather than the content itself, as shown in Fig. 1 (a). This bias has been prevalent in all types of LLM judges Qin et al. (2024); Li et al. (2023d), raising concerns about their reliability. Previous studies Zheng et al. (2024a, b); Zeng et al. (2023) have identified position bias alongside other biases and assessed its impact. Although mitigation strategies have been proposed, they often fall short due to incomplete bias elimination Guo et al. (2024), increased complexity Li et al. (2024b); Khan et al. (2024); Chua et al. (2024), the introduction of new biases Ohi et al. (2024), or inconsistent effectiveness Gallegos et al. (2024). These persistent challenges highlight the limitations of current approaches and point to a possible incomplete understanding of position bias. Although numerous studies have evaluated position bias in LLM-as-a-Judge settings, the underlying factors contributing to this phenomenon remain insufficiently explored.
In this study, we provide an in-depth and systematic investigation into position bias within the context of LLM-as-a-Judge. While evaluating Position Consistency, one of the most widely used methods for measuring position bias, we introduce two novel metrics: Preference Fairness and Repetition Stability. Specifically, we move beyond simply assessing Position Consistency by incorporating Preference Fairness, which provides deeper insights into the specific answer directions where models exhibit unfair preferences. Additionally, the measurement of Repetition Stability ensures that the observed position bias in the given model and tasks is not due to random variations, thus strengthening the reliability of the findings.
To investigate the underlying factors contributing to position bias, we categorized these factors into three levels: Judge-Level, Candidate-Level, and Task-Level. Our experiments are primarily conducted on pairwise comparisons, as LLM judges demonstrate superior performance in this setting. We further extend our study to more complicated list-wise comparison settings, involving evaluations of more than two candidate models by LLM judges. Our findings reveal several key insights: 1. The position bias of capable LLM judges is not a result of random variations. 2. There is a high volatility in the direction of preference, even within the same LLM judge when applied to different tasks. 3. Differences in answer quality among candidate models significantly influence position consistency. 4. Position bias is very weakly correlated with the length of prompts generated by the candidate models.
Building on these findings, we conduct an agreement analysis among the LLM judges. The results reveal that, although measures of position consistency may appear similar in general, judgments on specific instances vary significantly among LLM judges, even when they demonstrate comparable capabilities. Instances where numerous LLMs agree are generally easier to judge, whereas instances with disagreements are more challenging to evaluate and more prone to position bias. This analysis provides insights into the distribution of judging difficulty across the dataset and highlights the potential for dataset modifications by incorporating more instances that are either easier or more difficult to judge. Future work could explore how to measure the likelihood of position bias arise from the datasets by identifying and quantifying such hard-to-judge instances before implementing LLM judges.
2 Evaluation Settings & Definitions
We begin by outlining the settings for pairwise and list-wise comparisons employed in our experiments for LLM-as-a-Judge. Following this, we define the three metrics used in our evaluation: Position Consistency (PC), Preference Fairness (PF), and Repetition Stability (RS). Finally, we provide a detailed description of the factors we found that are related to position bias at the Judge-Level, Candidate-Level, and Task-Level.
2.1 Pairwise & List-wise Comparison
@@
Pairwise Comparison: In the context of pairwise comparison, LLM judges are tasked with selecting the better solution provided by two candidate models in response to a given task question. As shown in Fig. 1 (a), the system prompt, option choices, task question, and solutions from two candidate models (original prompt) are presented to the LLM judges to select the better solution. The experiment is conducted in a double-blind setting. The identities of the candidate models are hidden from the LLM judges, and the candidate models are unaware that their solutions will be compared to another model when answering the question. Then, the prompt with solutions in a swapped position (swapped prompt) is given to the same judge again, which results in a judgment pair. If the LLM judge consistently favors the same solution regardless of the swapped position, it is considered position consistent. Conversely, if the LLM judge selects different winners, position bias is observed, with the preference direction being either primacy (e.g. always choose $\{A\}$) or recency (e.g. always choose $\{B\}$). Example of measuring preference fairness with specific choice pairs is shown in Fig. 1 (b). To accommodate the possibility of ties, various option modes are employed: Two-Option mode restricts judges to choosing between two options, labeled A for the first candidate and B for the second. Three-Option mode introduces an additional choice, C, allowing judges to indicate a tie if neither solution is preferable, as illustrated in Fig. 1 (a). These option modes were explicitly specified in the system prompts to ensure clear guidance for the decision-making process of the LLM judges.
List-wise Comparison: Unlike pairwise settings, where LLM judges select the superior solution from two candidates, list-wise comparative approaches involve evaluating three or more candidates simultaneously, as shown in Fig. 1 (c). For efficiency, we prompt judges to select the best candidate rather than ranking the entire list. The “swapped setting" used in pairwise evaluations is generalized to order permutations for list-wise judgments, ensuring that each candidate appears in every possible position exactly once. For a list of $p$ candidates, this results in $p$ permutations. In the $i^{th}$ permutation, the $i^{th}$ candidate is set to appear in the first position. Additionally, an option is provided to account for ties, allowing judges to indicate if there is no certainly superior solution.
@@ 2.2 Evaluation Metrics
In our study, we first verify if the capable LLMs with a high repetition stability and then evaluate their position bias in terms of position consistency and preference fairness. The metrics are introduced as follows:
@@
Repetition Stability ($RS$) evaluates the reliability of LLM judges when presented with identical queries multiple times. It is essential to determine whether the judgments of LLMs, and consequently the observations of position bias, stem from a consistent evaluation pattern or by random variations. We measure this by calculating the percentage of the most frequent selections across multiple trials for each query, aggregated from all queries within each dataset. This metric is formalized as
$RC$ $=\frac{1}{N}\sum_{j=1}^{N}\frac{1}{n_{j}}\max_{k\in S}\left\{|C_{k}^{j}|\right\},$ (1)
@@
where $S=\{A,B,C,\dots\}$ refers to the set of choice options depending on the option mode, $|C_{k}^{j}|$ denotes the counts of each choice option selected by the judge for the $j^{th}$ query, $n_{j}$ represents the total number of repeating trials for that query, and $N$ is the total number of queries. The value of $RS$ ranges from a small positive value depending on the option mode, indicating completely random decisions, to 1.0, indicating perfect stability.
@@
Position Consistency ($PC$) quantifies how frequently a judge model prefers the same solution after the order of solutions is permuted. It is calculated as the ratio of consistent evaluation series to the total number of valid evaluations, where a series is deemed consistent if the judge model prefers the same winning solution across permutations. Formally, it is calculated as
$PC$ $=\frac{1}{n}\sum_{j=1}^{n}\mathbf{1}_{\{(C_{1}^{j},\dots,C_{p}^{j},)\in V\}},$ (2)
@@
where $V$ is the set of choices that correspond to position consistency, and $(C_{1}^{j},\dots,C_{p}^{j})$ denotes the judgment series for the $j^{th}$ query when there are $p$ candidate solutions in the list, and $n$ represents the number of prompt series. An example of such series of choices under pairwise comparison setting can be found in Fig. 1 (b). This formula aims to provide a direct measure of a judge model’s position bias and has been widely used in previous studies for its simplicity.
@@
Preference Fairness ($PF$) measures the extent to which judge models favor certain solution positions. In pairwise comparisons, an LLM judge may exhibit a preference for either primacy or recency. These terms replace the more verbose “preference for the first/second candidate model" used in previous studies, ensuring clarity and generalization for future research. The examples of such preferences are demonstrated in Fig. 1 (b). Previous studies proposed two common ways to measure the preference fairness. One way is to count the primacy-preferred and recency-preferred judgment pairs, which we termed as primacy-count-number ($pcn$) and recency-count-number ($rcn$). The counts are then normalized by the total number of prompt pairs (Zheng et al., 2024b; Zhu et al., 2023). However, the sensitivity of this measurement highly depends on the size of dataset, making comparisons across datasets unreliable, especially when the number of questions and instances varies for each task.
@@
Alternatively, instead of normalizing over the complete dataset, studies like (Li et al., 2023c; Liusie et al., 2024) treat position inconsistent evaluation instances independently. They calculate the percentages of primacy-preferred and recency-preferred judgment pairs relative to the total number of position inconsistent pairs. We denote these as inconsistent primacy rates ($ipr$) and inconsistent recency rates ($irr$), where $ipr+irr=1$. However, this approach overlooks the fact that “position consistent judgments are also preference fair”, which leads to overly penalizing highly consistent LLM-judges.
@@
$PF$ $=\frac{PF_{\text{raw}}-S^{-}_{\min}}{S^{+}_{\max}-S^{-}_{\min}}\times 2-1,$ (3)
$PF_{\text{raw}}$ $=(rcn\times irr)-(pcn\times ipr).$
where $S^{-}_{min}$ and $S^{+}_{max}$ are the minimum and maximum achievable $PF_{raw}$ scores for each judge on each task, respectively. This min-max scale ensures comparability across datasets by accounting for the range of achievable scores and centering the scale around zero. The $PF$ score is interpreted as follows:
@@
To extend this metric to list-wise comparisons, we employed a ‘one vs. all’ approach, defining primacy preference as favoring the first candidate solution while classifying all others as recency-preferred. This straightforward extension of the $PF$ computation maintains consistency with pairwise setups. By providing a single and comprehensive metric that applies to all evaluation instances and list-wise settings, our proposed $PF$ score ensures sensitivity across datasets, regardless of variations in the number of questions or instances, offering a significant improvement over previous methods.